Mathematics 376 – Probability and Statistics II First Lab/Discussion Group Assignments January 20, 2006

Group 1: Becca Bates, Luke Jaroszewicz, Michelle Hedrick Group 2: Natalie Cate, Jacqueline Varanelli, Courtney Walls Group 3: Gerry Bueti, Christina Roberto, Erin Cagianello Group 4: Angelina Kim, John Restivo, Sheena Stangler Group 5: Ingrid Gliottone, Erin Lord, Kaitlyn Mearn Group 6: Kaitlin Horndahl, Maggie Kielman, Tom Potvin Group 7: Kathryn Noonan, Lauren Pereira, Ben Daly Group 8: Packy Loughran, Jacqueline Markot, Katie McGovern, Beth Richard